cos^2-sin^2=2cos^2-1 prove that this is true
Answer by stanbon(75887)

cos^2-sin^2=2cos^2-1 prove that this is true
----------
Note: sin^2(x) = 1 - cos^2(x)
------------------------------------
Substitute that into the left side to get:
cos^2 -(1-cos^2) = 2cos^2-1
cos^2 -1 + cos^2 = 2cos^2-1
2cos^2 - 1 = 2cos^2 - 1
=============================
Cheers,
Stan H.